Determine minimum fluidization velocity of sand particle using water at 20 oC. The sand particles are mono-sized with surface volume spherical equivalent diameter dsv= 162 micron, a sphericity of 0.9

and density 2600 kg/m3


1
Expert's answer
2021-03-18T03:35:46-0400

{(150"\mu"umf) / ("\phi"s2dp2)} x {(1 - εmf) / εmf3)} + {(1.75"\rho"umf2) / ("\phi"sdp)} x {1 / εmf3)}

= g(ρh - "\rho")

dp = the diameter for spherical particles or the nominal diameter for irregular-shaped particles; ρ = the density of the fluid; ρh = the hudraulic density of the solid; g = the gravity acceleration constant, 9.81 m2/cm; μ = the dynamic viscosity of the fluid; εmf = the minimum fluidization voidage; umf = the minimum superficial fluidization velocity; and

ϕs = the sphericity of the particle.

dp = 162 micron = 162 x 10-6 m

"\rho"(water) = 998 kg/m3

"\rho"h(sand) = 2600 kg/m3

g = 9.81 m2/cm

μ = 1.002 x 10-3 Pa•s

εmf = (0.071 / ϕs)1/3 = (0.071 / 0.9)1/3 = 0.429.

{(150 x1.002 x umf) / (0.92(162 x 10-6)2} x {(1 - 0.429) / (0.429)3)} + {(1.7 x 998 x umf2) / (0.9 x 162 x 10-6)} x {1 / (0.429)3)}

= 9.81(2600 - 998)